# Efficient Manipulation-Enhanced Semantic Mapping With Uncertainty-Informed Action Selection

### v1

@@ IV-D2 Manipulation
To evaluate the influence of different uncertainty representations on our push sampling strategy, we conduct ablation experiments with two variants for distance map and pushing corridor generation: (i) using semantic uncertainty for distance map generation while using occupancy uncertainty for the pushing corridor ($\bf{u_{s}|u_{o}}$) to highlight semantically ambiguous but structurally visible regions, providing complementary cues for selecting informative push targets, the proposed method of this paper; and (ii) using occupancy uncertainty for both the uncertainty-based distance map and the pushing corridor generation ($\bf{u_{o}|u_{o}}$), We performed this ablation on the POMDP-CNABU-Informed push planner to have a more comparable evaluation that only focuses on the influence of the push selection, due to its systematic greedy search characteristics of the predefined fixed viewpoints.
The results after the final action step are presented in Tab. II. As seen from the last two rows, using $u_{s}$ as the base for distance map yields marginally higher semantic mIOU as it is better able to target regions with high semantic uncertainty. Furthermore, it leads to lower collision rates possibly due to clustering of semantically similar cells leading to pushes being executed on single objects reducing secondary pushes.
Our ablation studies on uncertainty metrics’ influence show that uncertainty is a good proxy for action selection between active sensor placement and manipulation. While semantic uncertainty can be used to target regions for active observation as well as for the push distance map, occupancy uncertainty can be effectively used to generate minimally invasive push actions.
@@ V Conclusion
In this paper, we presented a novel framework for efficient manipulation-enhanced semantic mapping that combines evidential metric-semantic map belief updates with uncertainty-aware action selection. By leveraging uncertainty estimates derived from the Beta and Dirichlet distributions, our system effectively selects informative next-best views through reinforcement learning and targets occlusion-critical objects for minimally invasive manipulations. Our integrated framework enables a robot to construct complete semantic maps in cluttered, occlusion-heavy shelf environments.
We demonstrated that our uncertainty-informed push planning approach outperforms the state-of-the-art method in terms of map accuracy, object displacement, and action efficiency. Furthermore, our approach is computationally efficient due to the reinforcement learning based viewpoint planning, and the informed push sampling strategy. These results highlight the importance of principled uncertainty-informed action selection for interactive perception in cluttered and confined spaces.

### v2

@@ IV-D2 Manipulation
To evaluate the influence of different uncertainty representations on our push sampling strategy, we conduct ablation experiments with two variants for distance map and pushing corridor generation: (i) using semantic uncertainty for distance map generation while using occupancy uncertainty for the pushing corridor ($\bf{u_{s}|u_{o}}$) to highlight semantically ambiguous but structurally visible regions, providing complementary cues for selecting informative push targets, the proposed method of this paper; and (ii) using occupancy uncertainty for both the uncertainty-based distance map and the pushing corridor generation ($\bf{u_{o}|u_{o}}$), We performed this ablation on the POMDP-CNABU-Informed push planner to have a more comparable evaluation that only focuses on the influence of the push selection, due to its systematic greedy search characteristics.
The results after the final action step are presented in Tab. IV. As seen from the last two rows, using $u_{s}$ as distance map yields marginally higher semantic mIOU as it is better able to target regions with high semantic uncertainty. Furthermore, it leads to lower collision rates possibly due to clustering of semantically similar cells leading to pushes being executed on single objects reducing secondary pushes.
Our ablation studies on the influence on the uncertainty metrics show that uncertainty is a good approximation for action selection between active sensor placement and manipulation. While semantic uncertainty can be used to target regions for active observation as well as for the push distance map, occupancy uncertainty can be effectively used to generate minimally invasive push actions.
@@ V Conclusion
In this paper, we presented a novel framework for efficient manipulation-enhanced semantic mapping that combines evidential metric-semantic map belief updates with uncertainty-aware action selection. By leveraging uncertainty estimates derived from the Beta and Dirichlet distributions for occupancy and semantics respectively, our system effectively selects informative next-best views through reinforcement learning and targets occlusion-critical objects for minimally invasive manipulations. As we showed in the experiments, out framework enables a robot to construct complete semantic maps in cluttered, occlusion-heavy shelf environments.
We demonstrated that our uncertainty-informed push planning approach outperforms the state-of-the-art method in terms of object displacement, action efficiency, and reduced number of fallen objects. It also shows comparable behavior in terms of map accuracy. Furthermore, our approach is highly computationally efficient due to the reinforcement learning based viewpoint planning and the informed push sampling strategy. These results highlight the importance of principled uncertainty-informed action selection for interactive perception in cluttered and confined spaces.
VI Acknowledgments
We would like to thank Joao Marcos Correia Marques and Murad Dawood for the fruitful discussions.